When your father was born 43 years ago, his grandparents deposited $500 in an account for him. Today, that account is worth $38,500. What was the annual rate of return on this account?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

38,500=500*(1+r/100)^43

(38,500/500)^(1/43)=(1+r/100)

(1+r/100)=1.1063

r=1.1063-1

=10.63%(Approx).